OpenStudy (aaalive):

Solve the equation: x+ 7/8= 4 (I need to find out what "x" is and I think it might be "3 1/8")

OpenStudy (anonymous):

Okay. We can see that x + 7/8 = 4. You can subtract 7/8 from both sides. Your answer is correct

OpenStudy (aaalive):

yay, thanks..do you mind helping with a few more?

OpenStudy (anonymous):

Yeah sure

OpenStudy (anonymous):

AAAlive are you there

OpenStudy (aaalive):

yeah, sorry my page glitched out

OpenStudy (aaalive):

I am just trying to take a screen shot of the question

OpenStudy (anonymous):

Okay

OpenStudy (aaalive):

I just took a screen shot of the question, how do I put it here?

OpenStudy (anonymous):

Hmm I dont know. Maybe click the button saying Attach File

OpenStudy (aaalive):

I don't think I have that option

OpenStudy (aaalive):

Ill just type it out the best I can

OpenStudy (anonymous):

Okay

OpenStudy (aaalive):

Solve. Express the answer in simplest form. -1/3 + w = 5 3/8 w =

OpenStudy (aaalive):

its a bit small...

OpenStudy (anonymous):

Yeah

OpenStudy (aaalive):

yeah...that link doesn't work

OpenStudy (aaalive):

but anyways... Solve. Express the answer in simplest form. -1/3 + w = 5 3/8 w =

OpenStudy (anonymous):

So we have -1/3 and w which equals 5 3/8. Now add -1/3 on each side

OpenStudy (aaalive):

I never got the "add "x" on each side" thing...especially with negative numbers

pooja195 (pooja195):

x+ 7/8= 4 Simply put, subtract 7/8 from both sides 4-7/8=?

OpenStudy (anonymous):

Its not x. Add -1/3

OpenStudy (aaalive):

I used x as to signify a random number...not specific to this question

pooja195 (pooja195):

3 1/8 is correct

pooja195 (pooja195):

You have it right

OpenStudy (aaalive):

I knew that...we're on a different question now

OpenStudy (anonymous):

AAAlive do you understand the step

OpenStudy (aaalive):

no, not really...if you explained it, sorry i didn't see it

pooja195 (pooja195):

It would be better if you reposted in on a new post.... -1/3 + w = 5 3/8 To make your life easier you can turn the mixed number into an improper fraction 43/8 then add 1/3 to cancel it out 43/8+1/3=?

OpenStudy (anonymous):

what pooja195 said is correct but one thing is wrong. She said 43/8 + 1/3 but you have to add negative 1/3 to 43/8 Then you will find out w

pooja195 (pooja195):

It's not wrong.

pooja195 (pooja195):

You would add 1/3 because you want to cancel out the negative.

OpenStudy (aaalive):

so...Im confused..do I add the negative or not

pooja195 (pooja195):